OpenStudy (anonymous):

What is the equation of a line perpendicular to 8x+9y=2 that passes through the point (7,-6)

OpenStudy (adamaero):

First, you need it in standard form y = mx + b

OpenStudy (adamaero):

Does the negative reciprocal of "m" sound familiar?

OpenStudy (anonymous):

Yes, is it 9/8?

OpenStudy (anonymous):

I think the answer might be Y=9/8x+2/9, but I'm not sure...

OpenStudy (adamaero):

use point slope form to use (7, -6)

OpenStudy (anonymous):

What is that?

OpenStudy (anonymous):

I see, so my answer is wrong?

OpenStudy (adamaero):

ya, I wish it was that easy tho man

OpenStudy (anonymous):

Me too, bro.

OpenStudy (anonymous):

Hey man, I just wanted to let you know that the right answer is Y=9/8x-111/8.